Required information

[The following information applies to the questions displayed below.]

The following data pertain to British Isles Aggregates Company, a producer of sand, gravel, and cement, for the year just ended.

Sales revenue £ 6,750,000
Cost of goods sold 2,235,000
Operating expenses 3,907,500
Average invested capital 1,687,500

£ denotes the British pound sterling, the national monetary unit of Great Britain.

  1. 3-a. Assuming that the expenses and cost of goods sold are reduced in order to improve the firm's ROI to 40 percent, compute the firm's new sales margin.

  2. 3-b. Show how the new sales margin and the old capital turnover together result in a new ROI of 40 percent.
Answer #1

3 a)Calculation of Sales Margin:-

Sales Revenue 6,750,000
Less: Cost of Goods Sold 2,235,000
Operating Expenses 3,907,500 6,142,500
Net Profit 607,500

Required ROI = 40%

Required Net Profit = (Sales Revenue*40)/100

=(1,687,500 *40)/100 = 675,000

New Sales Margin =(Net Profit/ Sales Revenue)*100

= (675,000/ 6,750,000)*100 = 10%

3 b) Old Capital Turnover Ratio = Sales / Average Capital Invested = 6,750,000/1,687,500 = 4 times

New Sales Margin = 10% (As Calculated in First Part)

ROI = New Sales Margin * Old Capital Turnover

ROI = 10%*4 times = 40%